# Background Semantics Matter: Cross-Task Feature Exchange Network for Clustered Infrared Small Target Detection With Sky-Annotated Dataset

### v1

@@ Abstract
Infrared small target detection poses unique challenges due to the scarcity of intrinsic target features and the abundance of similar background distractors. We argue that background semantics play a pivotal role in distinguishing visually similar objects for this task. To address this, we introduce a new task—-clustered infrared small target detection, and present DenseSIRST, a novel benchmark dataset that provides per-pixel semantic annotations for background regions, enabling the transition from sparse to dense target detection. Leveraging this dataset, we propose the Background-Aware Feature Exchange Network (BAFE-Net), which transforms the detection paradigm from a single task focused on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. BAFE-Net introduces a cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks. Furthermore, we propose the Background-Aware Gaussian Copy-Paste (BAG-CP) method, which selectively pastes small targets into sky regions during training, avoiding the creation of false alarm targets in complex non-sky backgrounds. Extensive experiments validate the effectiveness of BAG-CP and BAFE-Net in improving target detection accuracy while reducing false alarms. The DenseSIRST dataset, code, and trained models are available at https://github.com/GrokCV/BAFE-Net.
@@ I Introduction
Figure 1: The small target in the orange box is the real object we are interested in in the sky background, and the small target in the blue box is a virtual alarm in the background of the building. Figure (a) and Figure (c) show clustered small targets against the background of the sky and false alarms against the background of buildings, and figure (b) shows false alarms against the background of buildings.
Infrared imaging has emerged as a crucial tool in various domains, particularly in scenarios where long-distance imaging is required in the absence of visible light, such as night-time operations or environments with limited visibility [1]. By capturing the thermal distribution of objects, infrared imaging provides critical information for applications such as unmanned system monitoring [2], maritime target search [3], and rescue missions [4]. However, in these applications, the targets of interest are often located at great distances, resulting in their appearance as extremely small-sized objects in the captured images. Consequently, infrared small target detection plays a pivotal role in accurately interpreting and extracting valuable insights from these images, enabling the development of robust and efficient systems that can operate effectively in challenging conditions.
@@ I-A Prior Works on Clustered Infrared Small Target Detection
While the potential of infrared imaging has been recognized across various domains, existing infrared small target detection techniques are predominantly tailored for the detection of sparse targets [5, 6], which proves to be inadequate when faced with complex application scenarios such as drone swarm defense, multi-target tracking systems, and extensive maritime search and rescue operations. Consequently, there is an urgent need for advanced technology capable of effectively identifying and tracking dense infrared small targets.
As the distribution of small targets transitions from sparse to dense scenarios, existing infrared small target detection techniques face a multitude of challenges that drastically hinder their effectiveness. Firstly, methods that rely on the assumption of sparsely distributed targets become inapplicable in dense scenes, where targets are closely situated and may occlude one another, leading to the breakdown of single-target analysis strategies. Moreover, the reliance on local contrast features [7, 8], a staple of traditional approaches, becomes unreliable as the targets are no longer set against a uniform background but are instead interspersed with other proximate objects. Furthermore, even state-of-the-art deep learning methods [9, 10, 11, 12] struggle when confronted with dense target scenarios where the delineation between targets and background blurs, significantly adding to the complexity of detection tasks. Therefore, current methods are plagued by inefficiencies when operating in dense settings, frequently resulting in missed detections and high false alarm rates. This calls for new solutions tailored to overcome the intricacies of dense infrared small target detection.
From another perspective, despite the rapid strides made in generic object detection, with state-of-the-art methods like TOOD [13] and CrossKD [14] addressing dense object detection, these techniques cannot be directly transplanted to the domain of infrared small target detection. The challenges we face in this specialized field stem from several critical factors:
@@
    Abundance of Similar Background Distractors: Infrared imagery is often plagued by a multitude of background distractors that bear a striking resemblance to the targets of interest. These distractors, such as noise spots and hot spots, also manifest as blobs, making it nearly impossible to discriminate between them and genuine targets based on local visual comparisons.
These challenges inevitably lead us to ponder a fundamental question: What truly distinguishes visually similar objects in single-frame infrared small target detection, assigning them different class labels? Unraveling this enigma holds the key to developing effective and robust algorithms for infrared small target detection, enabling them to navigate the unique challenges posed by this task.
@@ I-B Motivation
However, it is crucial to highlight a significant gap in the current landscape of infrared small target detection research. Despite the availability of several public datasets, such as SIRST v1 [15] /v2 [4] , NUDT-SIRST [16], and IRSTD1K [12] , none of these datasets provide detailed semantic annotations for the background regions. To enable the explicit semantic modeling of backgrounds, a prerequisite is the creation of a dataset that includes background class annotations. Such a dataset would provide a solid foundation for researchers to develop algorithms that can leverage the rich contextual information present in the background regions.
@@
Building upon this foundation, we propose the Background-Aware Feature Exchange Network (BAFE-Net) framework, which transforms the infrared small target detection paradigm from a single target detection task that solely focuses on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. BAFE-Net introduces a cross-task feature hard-exchange module in the target detection head and background segmentation head, which embeds target semantics and background semantics between the segmentation and detection tasks through a feature adapter. To the best of our knowledge, this paper is the first to propose a cross-task feature hard-exchange mechanism. Benefiting from the explicit modeling of background semantics, BAFE-Net can better understand the contextual information in the scene, thereby significantly improving the accuracy of target detection and greatly reducing the false alarm rate.
@@
    We propose the BAFE-Net, which achieves a learning mechanism of mutual promotion between segmentation and detection tasks by implementing feature hard-exchange with a cross-task adapter, based on explicit background modeling.
To validate the effectiveness of BAG-CP and BAFE-Net, we provide extensive experimental results demonstrating their ability to improve the accuracy of target detection while reducing the false alarm rate. The experimental results demonstrate that our method outperforms the compared methods, showcasing its superior performance.
@@ II-A Infrared Small Target Detection
Infrared small target detection methods can be primarily categorized into model-driven and data-driven methods. Model-driven methods can be broadly categorized into three subclasses, each incorporating different strategies to model the relationship between targets and backgrounds.
@@
    Background Estimation Methods: These include techniques like Top-hat filtering [17] and Max-Median filtering [18]. By leveraging the local consistency of backgrounds, these methods construct filters to enhance target visibility. However, they falter when dealing with extremely weak targets, leading to missed detections and a high sensitivity to hyperparameters. The primary limitation stems from insufficient semantic understanding of the features, which often results in poor performance in complex backgrounds and increased false alarms.
@@
    Low-Rank and Sparse Decomposition (LRSD) Methods: This category includes approaches like the infrared patch-image model (IPI) [5], partial sum of tensor nuclear norm (PSTNN) [20], and sparse regularization-based twisted tensor model. These methods have gained popularity for their robustness in target and background separation. However, they severely suffer from the challenge of distinguishing background distractors from targets due to the similar sparse characteristic.
@@
Transitioning from model-driven methods, data-driven approaches have revolutionized the field by leveraging neural networks to automatically extract and learn features relevant for detecting small targets. Methods such as asymmetric context modulation (ACM) [15], dense nested attention networks (DNANet) [16], and gated shaped trans u-net (GSTUnet) [11] have shown significant promise. These methods concentrate on integrating the features of infrared small targets with neural networks, often utilizing deeper neural networks or more robust attention modules. The advantages of data-driven methods lie in their ability to automatically extract features and focus on multi-scale feature interaction, which can be seen as a form of implicit modeling of the relationship between targets and background.
While effective to a degree, these methods treat all backgrounds as a single class, ignoring the significant impact of different background semantics on distinguishing between true targets and false alarms with similar appearances. Moreover, despite not fully depending on sparse priors or local contrast priors like traditional methods, some approaches still leverage this prior knowledge to enhance detection performance. Their performance is greatly affected in dense scenes with multiple adjacent targets, where the assumption of local contrast between targets and their neighboring background no longer holds. This fundamental shift in target distribution necessitates the development of approaches that can effectively handle the complex interactions and dependencies among multiple targets in close proximity.
In contrast to previous studies, our work addresses several key limitations and explores new frontiers in infrared small target detection:
@@
    Focusing on the Unexplored Challenge of Clustered Targets: Unlike traditional approaches that primarily focus on isolated targets, we take on the more complex task of detecting clustered targets, introducing a new challenge in this field. We further facilitate the research in this area by providing DenseSIRST, a dedicated dataset created specifically for this novel task.
@@
    Incorporation of Explicit Semantic Modeling with DenseSIRST: In contrast to previous studies that often treat all backgrounds as a homogeneous class, we leverage the DenseSIRST dataset’s semantic segmentation annotations for background modeling. Our proposed BAFE-Net thereby gains an enhanced understanding of the target’s background, leading to improved detection accuracy.
@@
    Introduction of Cross-Task Feature Exchange: Our cross-task feature exchange module brings together background semantics and target features through a feature adapter, a strategy not commonly found in existing work. This approach enables a more precise differentiation between targets and their backgrounds.
@@ II-B Data Augmentation for Object Detection
Where our methodology aligns with the broader goal of enhancing infrared small target detection, it stands out in its unique focus and innovative strategies:
@@
    Pioneering Data Augmentation for Clustered Infrared Small Target Detection: Uniquely, our method pioneers a data augmentation technique specifically tailored for clustered infrared small target detection. It takes into account the local contrast prior and enables dynamic random variations in local contrast, a perspective not often considered in existing techniques.
@@
    Context-Aware Target Placement: Our proposed BAG-CP approach diverges from the commonly used random or fixed-position copy-paste tactics. Instead, it integrates background semantics to guide target placement, thereby skirting potential category conflicts induced by inconsistent background contextual semantics. This nuanced approach to the distinctive challenges in infrared small target detection tasks sets our work apart.
@@ II-C Multi-Task Learning and Cross-Task Mechanism
Multi-task learning (MTL) is a pivotal strategy in computer vision that enhances model performance and generalization by jointly learning related tasks, such as detection and segmentation. By learning shared representations from multiple supervisory tasks, MTL has shown remarkable success in the field of recognition.
Recent studies have introduced novel architectures and mechanisms to bolster the capabilities of MTL in object detection. Takikawa et al. [29] introduced a two-stream CNN architecture for semantic segmentation that explicitly wires shape information as a separate processing branch, using gates to connect the intermediate layers of the two streams. Misra et al. [30] proposed a new sharing unit called “cross-stitch” unit, which combines activations from multiple networks and can be trained end-to-end to learn an optimal combination of shared and task-specific representations. Zhang et al. [12] developed ISNet, which employs a Taylor finite difference (TFD)-inspired edge block and two-orientation attention aggregation (TOAA) block to enhance edge information and capture shape characteristics of infrared small targets.
However, these methods may not be well-suited for clustered infrared small target detection, as they often rely on edge detection or surface normal estimation, which assume larger targets. In real-world scenarios, infrared small targets typically consist of only 1-9 pixels and are subject to the diffusion effect, lacking clear edges and falling far below the conventional definitions of small targets, i.e., $32\times 32$ pixels in general object detection.
In this context, cross-task mechanisms that enable feature exchange between different tasks are more relevant to our work. These methods enhance MTL by facilitating feature exchange or distillation across different tasks, allowing each task to benefit from complementary signals of others. Methods like prediction-and-distillation network (PAD-Net) utilize attention-based mechanisms to distill multi-modal features effectively [31], while MTI-Net extends this concept with a multi-scale approach better suited for handling diverse task demands [32].
While our work builds upon the foundational principles of multi-task learning and cross-task mechanisms, it distinguishes itself in the feature hard-exchange mechanism. Diverging from the soft fusion approach used in attention-based methods, we employ a feature hard-exchange mechanism for feature fusion. This simple yet effective strategy ensures a more robust integration of features from different tasks.
@@ III The DenseSIRST Dataset
Figure 2: The four images above show diverse backgrounds, including city, sea, cloud, and mountain. The five diagrams below illustrate various targets, including Point, Cluster, UAV, Plane, and Ship.
Figure 3: A total of 12 dataset images are shown. In each set of images, the top figure shows the simulated dense small target, and the bottom figure shows the corresponding sky segmentation of the image.
@@ III-A Motivation and Rationale
The detection of small, low-contrast targets against complex backgrounds in infrared imagery is a significant challenge that impacts the performance of computer vision algorithms. While existing datasets have contributed to the advancement of this field, they often suffer from limitations such as sparse target distribution and a focus on larger target sizes, which do not adequately represent real-world scenarios.
To address these limitations, we introduce the DenseSIRST dataset, a richly annotated collection of infrared images featuring densely populated small targets. DenseSIRST not only increases the number of small targets per image but also encompasses a variety of target sizes and complex clustering, crucial for training and evaluating algorithms under realistic and challenging conditions.
TABLE I: Main characteristics of several popular SIRST datasets.
@@
SIRST [15] Real Pixel 427 533 23 Sparse
@@
Table I compares DenseSIRST with several well-regarded datasets, such as SIRST [15], SIRSTv2 [4], IRSTD1K [12], and SIRSTAUG [33]. These datasets predominantly feature sparsely distributed targets and lack comprehensive annotations for clustered small targets. In contrast, DenseSIRST distinguishes itself through:
  1. 1.
    Hybrid Image Composition: Combining real and synthetic targets, DenseSIRST enhances both the realism and diversity of the training data.
  2. 2.
    Target Density and Diversity: With an average target area significantly smaller than existing datasets and a higher number of targets per image, DenseSIRST provides a challenging scenario for detection algorithms.
  3. 3.
    Varied Background Scenarios: The dataset includes images from diverse backgrounds such as urban, mountainous, maritime, and cloudy scenes, as shown in Fig. 2, each contributing to the complexity of the detection tasks.
Fig. 3 showcases a selection of images from the DenseSIRST dataset, illustrating the dense distribution of small targets and the corresponding sky segmentation annotations. These annotations provide valuable contextual information for developing and evaluating detection algorithms.

### v2

@@ Abstract
Infrared small target detection poses unique challenges due to the scarcity of intrinsic target features and the abundance of similar background distractors. We argue that background semantics play a pivotal role in distinguishing visually similar objects for this task. To address this, we introduce a new task—-clustered infrared small target detection, and present DenseSIRST, a novel benchmark dataset that provides per-pixel semantic annotations for background regions, enabling the transition from sparse to dense target detection. Leveraging this dataset, we propose the Background-Aware Feature Exchange Network (BAFE-Net), which transforms the detection paradigm from a single task focused on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. BAFE-Net introduces a dynamic cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks. Furthermore, we propose the Background-Aware Gaussian Copy-Paste (BAG-CP) method, which selectively pastes small targets into sky regions during training, avoiding the creation of false alarm targets in complex non-sky backgrounds. Extensive experiments validate the effectiveness of BAG-CP and BAFE-Net in improving target detection accuracy while reducing false alarms. The DenseSIRST dataset, code, and trained models are available at https://github.com/GrokCV/BAFE-Net.
@@ I Introduction
Figure 1: Illustration of the background semantics’ crucial role in infrared small target detection. The small target in the orange box represents a genuine target of interest against the sky background, while the small target in the blue box is a false alarm in the context of the building background. When considering only local regions, it is challenging to differentiate between clustered small targets and false alarms due to their similar appearances. By incorporating global contextual information, the distinction between real targets and false alarms becomes more apparent.
Infrared imaging has emerged as a crucial tool in various domains, particularly in scenarios where long-distance imaging is required in the absence of visible light, such as night-time operations or environments with limited visibility [1]. By capturing the thermal distribution of objects, infrared imaging provides critical information for applications such as unmanned system monitoring, maritime target search [2], and rescue missions [3]. However, in these applications, the targets of interest are often located at great distances, resulting in their appearance as extremely small-sized objects in the captured images. Consequently, infrared small target detection plays a pivotal role in accurately interpreting and extracting valuable insights from these images.
@@ I-A Prior Works on Clustered Infrared Small Target Detection
While the potential of infrared imaging has been recognized across various domains, existing infrared small target detection techniques are predominantly tailored for the detection of sparse targets [4, 5], which proves to be inadequate when faced with complex application scenarios such as drone swarm defense, multi-target tracking systems, and extensive maritime search and rescue operations. Consequently, there is an urgent need for advanced technology capable of effectively identifying and tracking clustered infrared small targets.
As the distribution of small targets transitions from sparse to dense scenarios, existing infrared small target detection techniques face a multitude of challenges that drastically hinder their effectiveness. Firstly, methods that rely on the assumption of sparsely distributed targets become inapplicable in dense scenes, where targets are closely situated and may occlude one another, leading to the breakdown of single-target analysis strategies. Moreover, the reliance on local contrast features [6], a staple of traditional approaches, becomes unreliable as the targets are no longer set against a uniform background but are instead interspersed with other proximate objects. Furthermore, even state-of-the-art deep learning methods [7, 8, 9, 10] struggle when confronted with dense target scenarios where the delineation between targets and background blurs, significantly adding to the complexity of detection tasks. For instance, Dai et al. [11] introduced a detection transformer designed for arbitrarily oriented objects, yet its applicability to dense infrared small target settings remains limited due to the lack of intrinsic target features. Therefore, current methods are plagued by inefficiencies when operating in dense settings, frequently resulting in missed detections and high false alarm rates. This calls for new solutions tailored to overcome the intricacies of dense infrared small target detection.
While the rapid advancements in generic object detection have led to state-of-the-art methods like TOOD [12] and CrossKD [13] that address dense object detection, these techniques cannot be directly applied to the unique challenges posed by infrared small target detection. The specialized nature of this field presents several critical factors that hinder the direct transplantation of generic vision approaches:
@@
    Abundance of Similar Background Distractors: Infrared imagery is often plagued by a multitude of background distractors that bear a striking resemblance to the targets of interest. These distractors, such as noise spots and hot spots, also manifest as blobs, making it nearly impossible to discriminate between them and genuine targets based on local visual comparisons. Liao et al. [14] tackled similar challenges in RGB-thermal salient object detection, where distortions like noise and motion blur compounded the complexity of separating salient objects from background clutter.
These challenges inevitably lead us to ponder a fundamental question: What truly distinguishes visually similar objects in single-frame infrared small target detection, assigning them different class labels? Unraveling this enigma holds the key to developing effective and robust algorithms for this task.
@@ I-B Motivation
However, it is crucial to highlight a significant gap in the current landscape of infrared small target detection research. Despite the availability of several public datasets [15, 3, 16, 10], none of these datasets provide detailed semantic annotations for the background regions. To enable the explicit semantic modeling of backgrounds, a prerequisite is the creation of a dataset that includes background class annotations. Such a dataset would provide a solid foundation for researchers to develop algorithms that can leverage the rich contextual information present in the background regions.
@@
Building upon this foundation, we propose the Background-Aware Feature Exchange Network (BAFE-Net) framework, which transforms the infrared small target detection paradigm from a single target detection task that solely focuses on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. BAFE-Net introduces a dynamic cross-task feature hard-exchange module in the target detection head and background segmentation head, which selectively embeds target semantics and background semantics between the segmentation and detection tasks through a feature adapter. To the best of our knowledge, this paper is the first to propose a cross-task feature hard-exchange mechanism. Benefiting from the explicit modeling of background semantics, BAFE-Net can better understand the contextual information in the scene, thereby significantly improving the accuracy of target detection and greatly reducing the false alarm rate.
@@
    We propose the BAFE-Net, which achieves a dynamic mechanism of mutual promotion between segmentation and detection tasks by implementing selective feature hard-exchange with a cross-task adapter, based on explicit background modeling.
@@ II-A Infrared Small Target Detection
Infrared small target detection methods can be broadly classified into two main categories: model-driven and data-driven approaches. The model-driven methods, in particular, encompass three distinct subclasses, each employing unique strategies to capture and represent the relationship between targets and their surrounding backgrounds.
@@
    Background Estimation Methods: By leveraging the local consistency of backgrounds, these methods construct filters to enhance target saliency. However, they falter when dealing with extremely weak targets, leading to missed detections and a high sensitivity to hyperparameters. The limitation stems from insufficient semantic understanding of the features [17], which results in poor performance in complex backgrounds and increased false alarms.
@@
    Low-Rank and Sparse Decomposition (LRSD) Methods: This category includes approaches like the infrared patch-image (IPI) model [4] and reweighted infrared patch-tensor (RIPT) model [5]. These methods have gained popularity due to their superior background suppression capabilities compared to previous techniques. However, they severely suffer from the challenge of distinguishing background distractors from targets due to the similar sparse characteristic.
@@
Transitioning from model-driven methods, data-driven approaches have revolutionized the field by leveraging deep networks to automatically learn features relevant for detecting small targets. Recent advancements have also focused on implementing these deep learning-based methods efficiently, such as the work on high-accuracy and real-time two-stage small object detection on FPGA [19]. Methods such as asymmetric context modulation (ACM) [15], dense nested attention networks (DNANet) [16], and gated shaped trans U-Net (GSTUnet) [9] have shown significant promise. These methods concentrate on integrating the features of infrared small targets with neural networks, utilizing deeper neural networks or more robust attention modules. Similar advancements in applying deep learning to infrared imaging have been seen in related fields, such as the work on infrared-guided eye localization for autostereoscopic displays [20]. The advantages of data-driven methods lie in their ability to automatically extract features and focus on multi-scale feature interaction, which can be seen as a form of implicit modeling of the relationship between targets and background. Similar approaches have been applied in other domains, such as cloud detection, where multi-level information fusion with edge information injection has been used to improve detection performance [21]. In a related field, transformer-based models have been employed for spatio-temporal unsupervised traffic anomaly detection in aerial videos [22], demonstrating the versatility of these advanced deep learning techniques in handling complex visual data with small objects or anomalies.
While effective to a degree, these methods treat all backgrounds as a single class, ignoring the significant impact of different background semantics on distinguishing between true targets and false alarms with similar appearances. Moreover, despite not fully depending on sparse priors or local contrast priors like traditional methods, some approaches still leverage this prior knowledge to enhance detection performance [23]. Their performance is greatly affected in dense scenes with multiple adjacent targets, where the assumption of local contrast between targets and their neighboring background no longer holds. This fundamental shift in target distribution necessitates the development of approaches that can effectively handle the complex interactions and dependencies among multiple targets in close proximity.
In contrast to previous studies, our work explores a new detection paradigm for infrared small targets:
@@
    Unexplored Challenge for Clustered Targets: Unlike traditional approaches that focus on sparse targets, we take on the more complex task of detecting clustered targets, introducing a new challenge in this field. We further facilitate the research by providing DenseSIRST, a dedicated dataset created specifically for this novel task.
@@
    Explicit Semantic Modeling: In contrast to previous studies that treat backgrounds as a homogeneous class, we leverage the DenseSIRST dataset’s semantic annotations for background modeling. Our proposed BAFE-Net thereby gains an enhanced understanding of the target’s background, leading to improved detection accuracy.
@@
    Introduction of Cross-Task Feature Exchange: Our dynamic cross-task feature exchange module brings together background semantics and target features through a feature adapter, a strategy not found in existing work. This approach enables a more precise differentiation between targets and their backgrounds.
@@ II-B Data Augmentation for Object Detection
Although our BAG-CP aligns with the broader goal of enhancing infrared small target detection, it stands out in its unique focus and strategies:
@@
    Pioneering Data Augmentation for Clustered Infrared Small Target Detection: Uniquely, our method pioneers a data augmentation technique specifically tailored for clustered infrared small target detection. It takes into account the local contrast prior and enables dynamic random variations in local contrast, a perspective not considered in existing techniques.
@@
    Background-Aware Target Placement: Our BAG-CP approach diverges from the commonly used random or fixed-position copy-paste tactics. Instead, it integrates background semantics to guide target placement, thereby skirting potential category conflicts induced by inconsistent background contextual semantics.
@@ II-C Multi-Task Learning and Cross-Task Mechanism
Multi-task learning (MTL) is a pivotal strategy in computer vision that enhances model performance and generalization by jointly learning shared representations from related tasks, such as detection and segmentation. Takikawa et al. [32] introduced a two-stream CNN architecture for semantic segmentation that explicitly wires shape information as a separate processing branch, using gates to connect the intermediate layers of the two streams. Misra et al. [33] proposed a new sharing unit called “cross-stitch” unit, which combines activations from multiple networks and can be trained end-to-end to learn an optimal combination of shared and task-specific representations. Zhang et al. [10] developed ISNet, which employs a Taylor finite difference inspired edge block and two-orientation attention aggregation block to enhance edge information and capture shape characteristics of infrared small targets.
However, these methods may not be well-suited for clustered infrared small target detection, as they often rely on edge detection or surface normal estimation, which assume larger targets. In real-world scenarios, infrared small targets typically consist of only 1 to 9 pixels and are subject to the diffusion effect, lacking clear edges and falling far below the conventional definitions of small targets, i.e., $32\times 32$ pixels in general object detection.
In this context, cross-task mechanisms that enable semantic exchange between different tasks are more relevant to our work. These methods enhance MTL by facilitating feature exchange or distillation across different tasks, allowing each task to benefit from complementary signals of others. Methods like prediction-and-distillation network utilize attention-based mechanisms to distill multi-modal features [34], while MTI-Net extends this concept with a multi-scale approach better suited for handling diverse task demands [35].
While our research shares the underpinnings of multi-task learning and cross-task mechanisms, we pioneer a new approach to feature fusion. Instead of following the well-trodden path of soft fusion found in attention-based methods, our work introduces an innovative protocol–-a dynamic hard-exchange mechanism. To the best of our knowledge, this is the first work to explore the potential of dynamic hard-exchange of features between tasks.
@@ III The DenseSIRST Dataset
Figure 2: Illustration of selected images from our proposed DenseSIRST dataset. Each image pair consists of a simulated dense small target image (top) and its corresponding sky-segmented version (bottom). These images highlight the challenges posed by densely clustered targets within diverse backgrounds.
@@ III-A Motivation and Rationale
The detection of small, low-contrast targets against complex backgrounds in infrared imagery poses a significant challenge. Existing datasets, though instrumental in progressing the field, exhibit critical deficiencies. As shown in Table I, well-regarded datasets such as SIRST V1 [15], SIRST V2 [3], IRSTD1K [10], and SIRSTAUG [36] predominantly feature sparsely distributed targets and lack per-pixel background annotations. This stands in stark contrast to the dense, diminutive targets frequently encountered in real-world scenarios.
To bridge this gap, we present DenseSIRST: a meticulously annotated dataset of infrared images teeming with densely clustered, small targets, which is a more realistic and challenging benchmark for algorithm training and evaluation. Fig. 2 showcases a selection of images from the DenseSIRST dataset, illustrating the dense distribution of small targets and the corresponding sky segmentation annotations. These annotations provide valuable contextual information for developing and evaluating detection algorithms.
TABLE I: Comparison of the DenseSIRST dataset with other popular datasets used in infrared small target detection.
@@
SIRST V1 [15] Real Pixel 427 533 23 Sparse
@@
Figure 3: Statistical characteristics of our DenseSIRST dataset, underscoring the detection challenges it presents. (a) The distribution of target sizes, with the circle sizes in the scatter plot indicating the prevalence of each size category. (b) The distribution of local contrasts for small targets in the dataset, demonstrating that the dataset encompasses a broad range of contrasts, down to a minimum of 0.15. (c) The brightness distribution of small targets in the dataset, highlighting the wide spectrum from dark to light targets this dataset offers for detection tasks.